# Distillation-Enabled Knowledge Alignment Protocol for Semantic Communication in AI Agent Networks

### v1

@@ Abstract
Index Terms:
Semantic communications, knowledge alignment, knowledge distillation, low-rank adaptation.
@@ III-A Distillation of Expert Knowledge
where =$\mathcal{M}=\{1,...,M\}$. Given that $r$ is much smaller than $I_{m}$ and $O_{m}$, the parameter ratio (PR) between the DK and $\Delta\mathcal{P}$ can be very small, which is calculated by
@@
$\min_{\mathcal{D}}~{}J_{\mathrm{A}}(\mathcal{D})=\mathop{\mathbb{E}}\limits_{\bm{X}\sim\varGamma_{\bm{X}}}J\big{(}\bm{f}(\bm{X};\mathcal{P}_{\mathrm{pt}}\!+\!\mathcal{D}),\bm{f}(\bm{X};\mathcal{P}_{\mathrm{pt}}\!+\!\Delta\mathcal{P})\big{)}.$ (6)
@@
where $\alpha$ is the step size, $\mathcal{D}[l]$ denotes the DK of level $l\!\in\!\mathcal{L}$ with $\mathcal{D}\!=\!\mathcal{D}[L]$, and $l^{(t)}$ is randomly sampled in =$\mathcal{L}=\{1,\dots L\}$.
@@ IV-B Evaluation Results
We then evaluate the multi-level DK allocation. In Fig. 4(4(a)), we sample 100 networks and compare the average network losses of different algorithms. Optimal indicates solving the ILP using Gurobi [14]. Greedy is the proposed greedy algorithm. Genetic is the genetic algorithm (GA), a typical heuristic algorithm for integer programming [18]. Moreover, in the fully-store case, every agent stores the complete multi-level DK, which is used as the baseline. Fig. 4(4(a)) shows that the local optimum found by the greedy algorithm is very close to the global optimum, reducing the average network loss by 28% compared to the baseline. In contrast, the performance of GA deteriorates quickly as $N$ increases. In Fig. 4(4(b)), we compare the computation time required by different algorithms for $L=3$, $4$, and $5$. From the figure, the computation time grows exponentially in the optimal and genetic cases. In comparison, the proposed greedy algorithm remains highly efficient for $N=50$, significantly boosting the DK allocation.

### v2

@@ III-A Distillation of Expert Knowledge
where $\mathcal{M}=\{1,...,M\}$. Given that $r$ is much smaller than $I_{m}$ and $O_{m}$, the parameter ratio (PR) between the DK and $\Delta\mathcal{P}$ can be very small, which is calculated by
@@
$\min_{\mathcal{D}}~J_{\mathrm{A}}(\mathcal{D})=\hskip-3.00003pt\mathop{\mathbb{E}}\limits_{\bm{X}\sim\varGamma_{\bm{X}}}\hskip-3.00003ptJ\big(\bm{f}(\bm{X};\mathcal{P}_{\mathrm{pt}}\!+\!\mathcal{D}),\bm{f}(\bm{X};\mathcal{P}_{\mathrm{pt}}\!+\!\Delta\mathcal{P})\big).$ (6)
@@
where $\alpha$ is the step size, $\mathcal{D}[l]$ denotes the DK of level $l\!\in\!\mathcal{L}$ with $\mathcal{D}\!=\!\mathcal{D}[L]$, and $l^{(t)}$ is randomly sampled in $\mathcal{L}=\{1,\dots L\}$.
@@ IV-B Evaluation Results
We then evaluate the multi-level DK allocation. In Fig. 4(4(a)), we sample 100 networks and compare the average network losses of different algorithms. Optimal indicates solving the ILP using Gurobi [14]. Greedy is the proposed greedy algorithm. Genetic is the genetic algorithm (GA), a typical heuristic algorithm for integer programming [18]. Moreover, in the fully-store case, every agent stores the complete multi-level DK, which is used as the baseline. Fig. 4(4(a)) shows that the local optimum found by the greedy algorithm is very close to the global optimum, reducing the average network loss by 28% compared to the baseline. In contrast, the performance of GA deteriorates quickly as $N$ increases. In Fig. 4(4(b)), we compare the computation time required by different algorithms for $L=3$, 44, and $5$. From the figure, the computation time grows exponentially in the optimal and genetic cases. In comparison, the proposed greedy algorithm remains highly efficient for $N=50$, significantly boosting the DK allocation.